OpenStudy (anonymous):

log2(x-1)-log2(x+4)=-2

sam (.sam.):

Careful

sam (.sam.):

\[\log(a)-\log(b)\neq \log(ab)\]

OpenStudy (anonymous):

Oh that was a minus sign... Silly me

OpenStudy (anonymous):

\[\log_2(x-2)-\log_2(x+4)=\log_2\left(\frac{x-2}{x+4}\right)\] So you get \[\log_2\left(\frac{x-2}{x+4}\right)=-2\\ 2^{(\cdots)}=2^{-2}\\ \frac{x-2}{x+4}=\frac{1}{4}\]